If the integers m and n are chosen at random between 1 and 100, then the probability that the number of the form 7   + 7 is divisible by 5 equals

Grade:11

2 Answers

Aman Bansal
11 years ago

Dear Shiva,

7m + 7n = 7(m+ n). If this is divisible by 5, then m+n must be divisible by 5, since 7 is not. So the question is asking: if m and n are chosen randomly from between 1 and 100, what is the probability that m+n is divisible by 5.

Now, I assume the question means "If m and n are chosen randomly from between 1 and 100 inclusive"; the question should make clear whether 1 and 100 are permissible values of m and n. If the range is inclusive, then the answer will be 1/5. There are a few ways to see this; we could consider units digits, since we''re dealing with division by 5. No matter what I choose for m, there will be two possible units digits for n that will make m+n a multiple of 5. So for example, if m=23, then the units digit of n must either be 2 or 7 in order that m+n ends in 5 or 0. Since there are ten possible units digits for n in total, and two that make m+n a multiple of five, there is a 2/10 = 1/5 chance I choose a value of n that makes m+n divisible by 5.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

http://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

upender surepally
11 years ago

0.4